Diffusion Morning Report 01/28/10

Ted Berg tberg at stanford.edu
Thu Jan 28 06:31:11 PST 2010


Good Morning All,
    All gases are currently down after a problem in CISx. Hopefully 
issue will be resolved soon and gases restored. Looking into dep rate 
differences between Thermco Poly 1 and 2. Thermco LTO testing. Sem 
Hitachi has vacuum issues, FS here today. LTO going down tomorrow for 
pull and clean. No other issues at this time.

  	  	  	  	  	 
  	Date: 	1/28/10 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	up 	  	.257/30 	  	 
Tylan 8-SiGe 	yellow 	needs test 	7.45/30 	  	 
Tylan 9-Poly 	up 	  	7.36/30 	  	 
Tylan 10-Nitride 	up 	  	24.54/30 	  	 
Tylan 11-LTO 	up 	going down on Fri for pull and clean 	28.59/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	yellow 	needs test 	  	  	 
Thermco Poly2 	down 	checking SiH4 mfc 	  	  	 
Thermco Nitride1 	yellow 	uniformity issues ? 	  	  	 
Thermco bpsg 	future 	testing for qual 	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	A on line 412 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20100128/f5612dbe/attachment.html>


More information about the diffusion mailing list